# Towards Reliable Forgetting: A Survey on Machine Unlearning Verification, Challenges, and Future Directions

### v1

@@ 4. Parametric Unlearning Verification
Second-order forgetting methods have not only been widely explored in deep learning but have also made significant progress in graph neural networks (GNNs). Chien et al. (2022) introduce the first proven theoretical guarantee for approximate graph unlearning, focusing on addressing the issue of deletion requests for graph-structured data in GNNs. This method handles three types of unlearning requests: node feature unlearning, edge unlearning, and node unlearning. Through theoretical analysis of simple graph convolution and its generalized PageRank extension, the study provides a theoretical guarantee for approximate unlearning. Dong et al. (2024) effectively address unlearning requests by modeling the intermediate state of the optimization objective between instances with and without the data to be unlearned (such as nodes, edges, and attributes) and demonstrates a tighter bound in approximating actual GNN parameters. Recently, focusing on edge unlearning in graphs, Wu et al. (2023) introduce a new influence function that efficiently computes the required updates while accounting for the neighborhood of deleted edges. This method removes the requested edges from the GNN without requiring retraining, while also providing a certified unlearning guarantee. Nevertheless, these certified graph unlearning methods still require bounded model errors on precise node embeddings to maintain their certification guarantees. To address this challenge, Yi and Wei (2024) propose ScaleGUN, the first method to extend certified graph unlearning to billion-edge graphs. ScaleGUN integrates approximate graph propagation techniques into certified graph unlearning, offering certification guarantees for three unlearning scenarios: node feature, edge, and node unlearning.
4.2. Reproducibility Verification.
@@ 4.2.1. Proof-of-Unlearning (PoUL)
Thudi et al. (2022) propose a reproducible unlearning verification method inspired by the Proof-of-Learning (PoL) (Jia et al., 2021). PoL aims to verify whether a machine learning model, has undergone effective training. It does this by logging the training process, tracking model checkpoints, the data points used, and the hyperparameters associated with updates. These logs are typically represented as a series of tuples, such as $\{(w_{i},x_{i},g_{i})\}$, where $w_{i}$ represents the model weights, $x_{i}$ represents the data point, and $g_{i}$ represents the update rule. Auditors use these logs to reproduce checkpoints and compute the error between the actual and reproduced checkpoints. If the error is below a certain threshold, the update is considered valid, ensuring the correctness of the training process.
Expanding PoL to validate the correctness of ”forgetting” specific data points, Thudi et al. (2022) introduce the concept of Proof-of-Unlearning (PoUL), which ensures that certain data points are successfully excluded during retraining. PoUL uses PoL logs to prove that the data points marked for ”forgetting” were not used during model updates. Zhang et al. (2024a) identify two essential conditions for PoUL: reproducibility and removability.
@@
$\|w_{t}-g_{t}(w_{t-1},d_{t})\|\leq\epsilon$
where $w_{t}$ is the model parameter at step $t$, $g_{t}$ is the update function, $d_{t}$ is the data used for the step, and $\epsilon$ is the tolerated error.
@@
where $d_{t}$ is the dataset at step $t$, and $D_{f}$ is the set of data to be removed. Auditors verify the correctness of the ”forget” operation by checking whether PoUL meets these two conditions.
@@
To mitigate these risks, Weng et al. (2024) propose a solution based on Trusted Execution Environments (TEEs) tailored for SISA unlearning scenarios. This approach leverages Intel’s SGX (Software Guard Extensions) to ensure the integrity of data tracking and model training. The solution consists of two layers: the authentication layer, which tracks the impact of data on submodels and updates them accordingly after data deletion, and the proof layer, which guarantees the correctness of both the learning and inference processes by verifying that the designated data has been properly removed.
Additionally, Eisenhofer et al. (2022) introduce a cryptographic verifiable machine unlearning method that employs SNARKs and hash chains for verifiable computation and proof generation. This framework is compatible with various unlearning techniques and has been validated through linear regression, logistic regression, and neural networks. The protocol ensures that users can verify the correctness of the unlearning process while maintaining model security and integrity.
@@ 5. Advantages and Limitations
  • Q1: Theoretical Guarantee: Does the method provide provable guarantees of forgetting?
  • Q2: Access Requirement: Can the method operate in a black-box setting, where internal model parameters are inaccessible?
  • Q3: Sample-Level Verification: Can the method evaluate forgetting at the level of individual samples?
  • Q4: Verification Accuracy: Does the method offer high verification accuracy?
  • Q5: No Pre-Injected Data Required: Is the method effective without requiring specially crafted or pre-injected data during training?
  • Q6: Efficiency and Scalability: Is the method computationally efficient and scalable to large models?
  • Q7: Method Specificity: Is the method general-purpose, or is it tailored to specific types of unlearning algorithms?
@@ 5.2.1. Fingerprint Verification
Fingerprint verification methods, including adversarial example fingerprints, backdoor fingerprints, and dataset watermarking.
Q1: Theoretical Guarantee and Q2: Access Requirement. All existing fingerprint verification methods lack formal theoretical guarantees regarding the effectiveness of the unlearning process. Consequently, they are inherently empirical in nature, and their results can only serve as indicative evidence rather than providing definitive assurance (For All: Q1: ). As for access requirements, all existing passive fingerprint methods require white-box access (Q2: ). In contrast, most active fingerprint methods support black-box verification, with the exception of Tam et al. (2024), which operates under a white-box setting (Q2: ).
Q3: Sample-Level Verification. Active fingerprint verification typically evaluates forgetting effectiveness by analyzing the collective activation patterns of a group of embedded samples, which makes it difficult to determine whether a specific data point has been successfully unlearned. To address this limitation, recent work (Xu et al., 2024b) introduces a backdoor-based verification mechanism that constructs dedicated sample pairs for each target data point, thereby enabling sample-level unlearning verification. Although this approach provides fine-grained verification capabilities, it relies on uploading a substantial amount of non-forgettable data, which may complicate its integration into privacy-sensitive or constrained application settings ().
@@
Q4: Verification Accuracy. The accuracy of fingerprint verification methods largely depends on factors such as the quality of the injected fingerprints and the specifics of the model’s training process (For All: Q4: ). For example, active fingerprint methods can be sensitive to changes in model architecture or security defenses (e.g., backdoor removal techniques). In contrast, passive fingerprint methods often rely on adversarial examples, whose effectiveness may significantly degrade in the presence of adversarial defense strategies, thereby undermining the reliability of the verification.
Q5: No Pre-Injected Data Required. Active fingerprint methods rely heavily on pre-injected data (Q5: ). This means that special samples must be embedded during training to enable subsequent verification, which limits their applicability in scenarios where post-hoc verification on arbitrary user data is required. In contrast, passive fingerprint methods offer greater flexibility, as they typically support verification through post-training queries without requiring any intervention during the training process (Q5: ).
Q6: Efficiency and Scalability and Q7: Method Specificity. Fingerprint verification methods are generally efficient and highly scalable. Since they do not require retraining the model, these methods can be deployed quickly and are applicable to large-scale models and datasets (For All: Q6: ). Moreover, fingerprint does not depend on any specific unlearning strategy, making it broadly compatible with a wide range of unlearning approaches. This generality enables fingerprint methods to maintain strong adaptability and practical utility across diverse application scenarios (For All: Q7: ).
@@ 5.2.2. Privacy Analysis Verification
Privacy analysis verification methods can be broadly categorized into data inference and data reconstruction approaches. Q1: Theoretical Guarantee and Q2: Access Requirement. Privacy analysis verification methods lack formal theoretical guarantees for the unlearning process, as their validation relies primarily on empirical evaluations via privacy analysis and does not offer rigorous theoretical support (Q1: ).
In terms of access requirements, privacy analysis methods exhibit a range of access assumptions. For example, data reconstruction techniques vary in their requirements, with some relying on black-box access (Graves et al., 2021; Xu et al., 2024a) and others on white-box access  (Bertran et al., 2024; Hu et al., 2024)(Q2: ). Mainstream data inference methods also frequently operate under black-box settings, although recent studies (Wang et al., 2024a) propose white-box variants to improve inference accuracy, thus they are rated as in Q2.
Q3: Sample-Level Verification and Q4: Verification Accuracy. Data inference methods enable sample-level verification by assessing whether the target model retains privacy-relevant information associated with specific data samples (Q3: ). While most data reconstruction methods are capable of recovering information at the sample level, the study (Graves et al., 2021) only supports class-level reconstruction and lacks the granularity required for single-sample verification (Q3: ).
In addition, the accuracy of these methods largely depends on factors such as model architecture, training dynamics, and the effectiveness of the attack design. For instance, the quality of the attack model, shifts in the training data distribution, and model updates can all affect the stability of the verification results. Due to these uncertainties, the verification accuracy of privacy analysis methods is generally unstable (Q4: ).
Q5: No Pre-Injected Data Required, Q6: Efficiency and Scalability, and Q7: Method Specificity. Privacy analysis verification methods generally do not require pre-injected special samples (Q5:), making them more flexible and applicable for post-hoc evaluations. However, computational efficiency varies depending on the attack type used. Data reconstruction methods do not require retraining the model and rely on optimization or analytical techniques, making them relatively lightweight with good efficiency and scalability (Q6). In contrast, data inference methods usually require training multiple auxiliary models, which incurs significant computational overhead in large-scale applications (Q6:). Additionally, privacy analysis methods are broadly applicable and can be used to assess various types of unlearning processes, so they score on Q7.
@@
Q1: Theoretical Guarantee and Q2: Access Requirement model performance serves as an intuitive indicator but does not offer formal theoretical guarantees for the unlearning process, and is thus rated as (Q1: ). In terms of access requirements, methods based on accuracy and relearn time typically require only black-box access, without the need for internal model parameters or training data, which greatly enhances their flexibility and applicability (Q2: ). Output methods also mostly relies on black-box outputs, although some studies (Vidal et al., 2024; Jeon et al., 2024b) suggest that using intermediate layer outputs can yield more effective verification results (Q2: ). In contrast, model sensitivity methods depend on a white-box setting, as they require access to model parameters to compute gradients (Q2: ).
Q3: Sample-Level Verification and Q4: Verification Accuracy Methods based on task quality and relearn time generally focus on overall performance on the unlearned dataset and cannot directly evaluate individual samples (Q3: ). In comparison, model output and model sensitivity methods assess the model’s response to specific samples and can thus provide more fine-grained, sample-level verification (Q3: ). However, since these methods rely on model behavior, they are sensitive to factors such as data distribution shifts and hyperparameter tuning, leading to performance fluctuations and making their verification accuracy and stability unreliable (Q4: ).
Q5: No Pre-Injected Data Required, Q6: Efficiency and Scalability, and Q7: Method Specificity Model performance methods do not rely on pre-injected data for verification (Q5: ). Methods based on relearn time, model output, and model sensitivity are efficient and scalable, as they do not require retraining the entire model. This makes them suitable for large models and datasets (Q6: ). For task quality , most approaches allow efficient evaluation by measuring how well the model completes the task. Recently, Nguyen et al. (2022b) propose a method that retrains a separate model and compares its task quality with that of the unlearned model, which results in significant computational overhead (Q6: ). As for method specificity, performance methods apply to various unlearning mechanisms (Q7: ).
@@ 5.3.1. Differential Privacy Verification
Q1: Theoretical Guarantee. Differential privacy verification methods are supported by a strong theoretical foundation, based on rigorous mathematical definitions. These methods provide quantifiable guarantees for the forgetting process through mathematical proofs and privacy protection algorithms (Q1: ).
Q2: Access Requirement. Differential privacy methods typically rely on white-box access to the model’s internal parameters. This is because to verify whether the model satisfies differential privacy conditions, auditors must have access to the model’s parameters (Q2: ).
Q3: Sample-Level Verification and Q4: Verification Accuracy. Differential privacy methods offer fine-grained control over each data point, enabling sample-level verification (Q3: ). Additionally, since they are based on solid theoretical guarantees, these methods are more trustworthy than behavioral approaches (Q4: ).
Q5: No Pre-Injected Data Required, Q6: Efficiency and Scalability and Q7: Method Specificity. Differential privacy methods do not rely on pre-injected specific samples; instead, the verification process only requires the use of existing data (Q5: ). While these methods offer strong theoretical guarantees, their computational efficiency and scalability are limited. Specifically, first-order methods necessitate the verification of each model parameter and gradient, whereas second-order methods involve costly Hessian matrix calculations, resulting in significant computational overhead. Additionally, most of these methods rely on strong assumptions, such as convexity or strong convexity, which further restrict their scalability (Q6: ).
Moreover, differential privacy methods can only enable effective unlearning verification when used in conjunction with their corresponding training mechanisms or algorithmic frameworks, which limits their applicability in certain unlearning scenarios (Q7: ).

### v2

@@ 4.1. Differential Privacy Verification
Second-order forgetting methods have not only been widely explored in deep learning but have also made significant progress in graph neural networks (GNNs). Chien et al. (2022) introduced the first proven theoretical guarantee for approximate graph unlearning, focusing on addressing the issue of deletion requests for graph-structured data in GNNs. This method handles three types of unlearning requests: node feature unlearning, edge unlearning, and node unlearning. Through theoretical analysis of simple graph convolution and its generalized PageRank extension, the study provides a theoretical guarantee for approximate unlearning. Dong et al. (2024) effectively addressed unlearning requests by modeling the intermediate state of the optimization objective between instances with and without the data to be unlearned (such as nodes, edges, and attributes) and demonstrates a tighter bound in approximating actual GNN parameters. Recently, focusing on edge unlearning in graphs, Wu et al. (2023) introduced a new influence function that efficiently computes the required updates while accounting for the neighborhood of deleted edges. This method removes the requested edges from the GNN without requiring retraining, while also providing a certified unlearning guarantee. Nevertheless, these certified graph unlearning methods still require bounded model errors on precise node embeddings to maintain their certification guarantees. To address this challenge, Yi and Wei (2024) proposed ScaleGUN, the first method to extend certified graph unlearning to billion-edge graphs. ScaleGUN integrates approximate graph propagation techniques into certified graph unlearning, offering certification guarantees for three unlearning scenarios: node feature, edge, and node unlearning.
4.2. Reproducibility Verification
@@ 4.2.1. Proof-of-Unlearning (PoUL)
Thudi et al. (2022) proposed a reproducible unlearning verification method inspired by the Proof-of-Learning (PoL) (Jia et al., 2021). PoL aims to verify whether a machine learning model has undergone effective training. It does this by logging the training process, tracking model checkpoints, the data points used, and the hyperparameters associated with updates. These logs are typically represented as a series of tuples, such as $\{(\mathbf{w}_{i},\mathbf{x}_{i},g_{i})\}$, where $\mathbf{w}_{i}$ represents the model weights, $\mathbf{x}_{i}$ represents the data point, and $g_{i}$ represents the update rule. Auditors use these logs to reproduce checkpoints and compute the error between the actual and reproduced checkpoints. If the error is below a certain threshold, the update is considered valid, ensuring the correctness of the training process.
Expanding PoL to validate the correctness of ”forgetting” specific data points, Thudi et al. (2022) introduced the concept of PoUL, which ensures that certain data points are successfully excluded during retraining. PoUL uses PoL logs to prove that the data points marked for ”forgetting” were not used during model updates. Zhang et al. (2024a) identified two essential conditions for PoUL: reproducibility and removability.
@@
$\|\mathbf{w}_{t}-g_{t}(\mathbf{w}_{t-1},d_{t})\|\leq\epsilon$
where $\mathbf{w}_{t}$ is the model parameter at step $t$, $g_{t}$ is the update function, $d_{t}$ is the data used for the step, and $\epsilon$ is the tolerated error.
@@
where $d_{t}$ is the dataset at step $t$, and $D_{f}$ is the set of data to be removed. Auditors verify the correctness of the ”forget” operation by checking whether PoUL meets these two conditions. These methods required recording and storing model parameters at every update step, which introduced significant storage overhead. To address this issue, some studies focused only on the final step of the unlearning process. For example, Wu et al. (2020) determined whether a predefined structural constraint was satisfied by comparing the parameters of a model retrained on $D_{r}$ with those of the target model. Nguyen et al. (2022b) further improved the distance metric by using the Frobenius norm of the difference between the corresponding weight matrices as a quantitative measure. These approaches did not rely on additional empirical behavior-level observation modules, but instead provided a formal verification pathway by directly measuring parameter-level deviations.
@@
To mitigate these risks, Weng et al. (2024) proposed a solution based on Trusted Execution Environments (TEEs) tailored for SISA unlearning scenarios. This approach leverages Intel’s Software Guard Extensions (SGX) to ensure the integrity of data tracking and model training. The solution consists of two layers: the authentication layer, which tracks the impact of data on submodels and updates them accordingly after data deletion, and the proof layer, which guarantees the correctness of both the learning and inference processes by verifying that the designated data has been properly removed.
Additionally, Eisenhofer et al. (2025) introduced a cryptographic verifiable machine unlearning method that employs SNARKs and hash chains for verifiable computation and proof generation. This framework is compatible with various unlearning techniques and has been validated through linear regression, logistic regression, and neural networks. The protocol ensures that users can verify the correctness of the unlearning process while maintaining model security and integrity.
@@ 5. Advantages and Limitations
  • Q1. Theoretical Guarantee: Does the method provide formal, provable guarantees of forgetting?
  • Q2. Access Requirement: Can the method operate in a black-box setting, where internal model parameters are inaccessible?
  • Q3. Sample-Level Verification: Can the method evaluate forgetting at the level of individual samples?
  • Q4. Verification Accuracy: Does the method offer high verification accuracy?
  • Q5. No Pre-Injected Data Required: Is the method effective without requiring specially crafted or pre-injected data during training?
  • Q6. Efficiency and Scalability: Is the method computationally efficient and scalable to large models?
  • Q7. Method Specificity: Is the method general-purpose, or is it tailored to specific types of unlearning algorithms?
@@ 5.2.1. Fingerprint Verification
Fingerprint verification methods, including passive fingerprint and active fingerprint.
Q1 Theoretical Guarantee and Q2 Access Requirement. All existing fingerprint verification methods lack formal theoretical guarantees regarding the effectiveness of the unlearning process. Consequently, they are inherently empirical in nature, and their results can only serve as indicative evidence rather than providing definitive assurance (For All: Q1. ). As for access requirements, all existing passive fingerprint methods require white-box access (Q2. ). In contrast, most active fingerprint methods support black-box verification, with the exception of Tam et al. (2024), which operates under a white-box setting (Q2. ).
Q3 Sample-Level Verification. Active fingerprint verification typically evaluates forgetting effectiveness by analyzing the collective activation patterns of a group of embedded samples, which makes it difficult to determine whether a specific data point has been successfully unlearned. To address this limitation, Xu et al. (2024b) introduced a backdoor-based verification mechanism that constructs dedicated sample pairs for each target data point, thereby enabling sample-level unlearning verification. Although this approach provides fine-grained verification capabilities, it relies on uploading a substantial amount of non-forgettable data, which may complicate its integration into privacy-sensitive or constrained application settings ().
@@
Q4 Verification Accuracy. The accuracy of fingerprint verification methods largely depends on factors such as the quality of the injected fingerprints and the specifics of the model’s training process (For All: Q4. ). For example, active fingerprint methods can be sensitive to changes in model architecture or security defenses (e.g., backdoor removal techniques). In contrast, passive fingerprint methods often rely on adversarial examples, whose effectiveness may significantly degrade in the presence of adversarial defense strategies, thereby undermining the reliability of the verification.
Q5 No Pre-Injected Data Required. Active fingerprint methods rely heavily on pre-injected data (Q5. ). This means that special samples must be embedded during training to enable subsequent verification, which limits their applicability in scenarios where post-hoc verification on arbitrary user data is required. In contrast, passive fingerprint methods offer greater flexibility, as they typically support verification through post-training queries without requiring any intervention during the training process (Q5. ).
Q6 Efficiency and Scalability and Q7 Method Specificity. Fingerprint verification methods are generally efficient and highly scalable. Since they do not require retraining the model, these methods can be deployed quickly and are applicable to large-scale models and datasets (For All: Q6. ). Moreover, fingerprint does not depend on any specific unlearning strategy, making it broadly compatible with a wide range of unlearning approaches. This generality enables fingerprint methods to maintain strong adaptability and practical utility across diverse application scenarios (For All: Q7. ).
@@ 5.2.2. Privacy Analysis Verification
Privacy analysis verification methods can be broadly categorized into data inference and data reconstruction approaches.
Q1 Theoretical Guarantee and Q2 Access Requirement. Privacy analysis verification methods lack formal theoretical guarantees for the unlearning process, as their validation relies primarily on empirical evaluations via privacy analysis and does not offer rigorous theoretical support (Q1. ).
In terms of access requirements, privacy analysis methods exhibit a range of access assumptions. For example, data reconstruction techniques vary in their requirements, with some relying on black-box access (Graves et al., 2021; Xu et al., 2024a) and others on white-box access (Bertran et al., 2024; Hu et al., 2024) (Q2. ). Mainstream data inference methods also frequently operate under black-box settings, although recent studies (Wang et al., 2024a) proposed white-box variants to improve inference accuracy, thus they are rated as in Q2.
Q3 Sample-Level Verification and Q4 Verification Accuracy. Data inference methods enable sample-level verification by assessing whether the target model retains privacy-relevant information associated with specific data samples (Q3. ). While most data reconstruction methods are capable of recovering information at the sample level, the study (Graves et al., 2021) only supports class-level reconstruction and lacks the granularity required for single-sample verification (Q3. ).
In addition, the accuracy of these methods largely depends on factors such as model architecture, training dynamics, and the effectiveness of the attack design. For instance, the quality of the attack model, shifts in the training data distribution, and model updates can all affect the stability of the verification results. Due to these uncertainties, the verification accuracy of privacy analysis methods is generally unstable (Q4. ).
Q5 No Pre-Injected Data Required, Q6 Efficiency and Scalability, and Q7 Method Specificity. Privacy analysis verification methods generally do not require pre-injected special samples (Q5. ), making them more flexible and applicable for post-hoc evaluations. However, computational efficiency varies depending on the attack type used. Data reconstruction methods do not require retraining the model and rely on optimization or analytical techniques, making them relatively lightweight with good efficiency and scalability (Q6. ). In contrast, many data inference methods typically require training multiple auxiliary models, which can incur significant computational overhead in large-scale applications. Recently, Wang et al. (2024a) demonstrated that inference verification can be performed without relying on auxiliary models by directly leveraging the unlearned model itself, thereby improving efficiency to some extent (Q6. ). Additionally, privacy analysis methods are broadly applicable and can be used to assess various types of unlearning processes, so they score on Q7.
@@
Q1 Theoretical Guarantee and Q2 Access Requirement. Model performance serves as an intuitive indicator but does not offer formal theoretical guarantees for the unlearning process, and is thus rated as (Q1. ). In terms of access requirements, methods based on accuracy and relearn time typically require only black-box access, without the need for internal model parameters or training data, which greatly enhances their flexibility and applicability (Q2. ). Output methods also mostly relies on black-box outputs, although some studies (Vidal et al., 2024; Jeon et al., 2024) suggested that using intermediate layer outputs can yield more effective verification results (Q2. ). In contrast, model sensitivity methods depend on a white-box setting, as they require access to model parameters to compute gradients (Q2. ).
Q3 Sample-Level Verification and Q4 Verification Accuracy. Methods based on task quality and relearn time generally focus on overall performance on the unlearned dataset and cannot directly evaluate individual samples (Q3. ). In comparison, model output and model sensitivity methods assess the model’s response to specific samples and can thus provide more fine-grained, sample-level verification (Q3. ). However, since these methods rely on model behavior, they are sensitive to factors such as data distribution shifts and hyperparameter tuning, leading to performance fluctuations and making their verification accuracy and stability unreliable (Q4. ).
Q5 No Pre-Injected Data Required, Q6 Efficiency and Scalability, and Q7 Method Specificity. Model performance methods do not rely on pre-injected data for verification (Q5. ). Methods based on relearn time, model output, and model sensitivity are efficient and scalable, as they do not require retraining the entire model. This makes them suitable for large models and datasets (Q6. ). For task quality, most approaches allow efficient evaluation by measuring how well the model completes the task. Recently, Nguyen et al. (2022b) proposed a method that retrains a separate model and compares its task quality with that of the unlearned model, which results in significant computational overhead (Q6. ). As for method specificity, performance methods apply to various unlearning mechanisms (Q7. ).
@@ 5.3.1. Differential Privacy Verification
Q1 Theoretical Guarantee. Differential privacy verification methods are supported by a strong theoretical foundation, based on rigorous mathematical definitions. These methods provide quantifiable guarantees for the forgetting process through mathematical proofs and privacy protection algorithms (Q1. ).
Q2 Access Requirement. Differential privacy methods typically rely on white-box access to the model’s internal parameters. This is because to verify whether the model satisfies differential privacy conditions, auditors must have access to the model’s parameters (Q2. ).
Q3 Sample-Level Verification and Q4 Verification Accuracy. Differential privacy methods offer fine-grained control over each data point, enabling sample-level verification (Q3. ). Additionally, since they are based on solid theoretical guarantees, these methods are more trustworthy than behavioral approaches (Q4. ).
Q5 No Pre-Injected Data Required, Q6 Efficiency and Scalability and Q7 Method Specificity. Differential privacy methods do not rely on pre-injected specific samples; instead, the verification process only requires the use of existing data (Q5. ). While these methods offer strong theoretical guarantees, their computational efficiency and scalability are limited. Specifically, first-order methods necessitate the verification of each model parameter and gradient, whereas second-order methods involve costly Hessian matrix calculations, resulting in significant computational overhead. Additionally, most of these methods rely on strong assumptions, such as convexity or strong convexity, which further restrict their scalability (Q6. ).
Moreover, differential privacy methods can only enable effective unlearning verification when used in conjunction with their corresponding training mechanisms or algorithmic frameworks, which limits their applicability in certain unlearning scenarios (Q7. ).